# On Steerability Factors for Growing Vine Robots

### v1

@@ IV-C Experimental Setup
We conducted a series of planar bending trials with both body types (Fig. 5). The robots were held at a constant length of 1.75 m using a motorized spool within a vine robot base similar to the one in [4]. Pressure was controlled using the same pressure regulators as in Sec. III-B. For each trial, a single actuator was aligned with the measurement plane, isolating its contribution to steering. Unlike the free-space trials in Sec. III, these experiments were performed with the robot lying on the floor, which reduced the influence of body weight, length, and gravitational loading along the body, allowing more direct characterization of fabrication methods and pressure ratio alone.
@@
The robot body was tracked using motion capture LED markers with the same system as in Sec. III-B, placed along the centerline at 25 cm spacing, starting at the base. This enabled reconstruction of the planar backbone for consistent extraction of the tip position and curvature. Similar to the previous experiments, we also placed LED markers aligned with the desired coordinate frame–this time with $y$ horizontal along the robot’s axis from base to tip when straight and $z$ up–and an LED at the base of the robot.
@@ V-B1 Diameter
The diameter of a vine robot should be chosen such that it will not collapse under the desired operating tip load, pressure, and length in free space. Its minimal impact on steerability suggests it should be largely based on the needs of the given environment.
@@ V-B2 Fabrication Method
Whether the exterior pouch fabrication or the integrated pouch fabrication should be chosen for a particular task depends on at what pressure ratio the task will be able to be completed. For tasks that must be completed at high pressure ratios (usually due to needing a high chamber pressure to achieve growth or prevent robot collapse), the exterior pouch design should yield higher curvature. For tasks that can be completed at high pressure ratios (i.e., ones where the robot body is supported and/or not actively growing while steering), the integrated pouch design should yield higher curvature.
@@ V-B3 Length
As long as the robot is not long enough to collapse, the robot’s free length should be maximized. While the free length of the robot during exploration of a space depends largely on the shape of the space being explored, an informed operator can choose which path to control the robot to take. For example, given multiple options of apertures to travel through before executing a sharp turn, an operator might choose the aperture that happens earlier in the growth process so that the robot is left with a longer length to curve to achieve the turn.
@@ V-B4 Tip Load
The total weight of the sensor payload on the robot’s tip should be as light as possible to minimize the impact on steerability. However, making certain sensors removable may also aid the robot in performing more difficult maneuvers in situations where not all sensors are needed.
@@ V-B5 Operating Pressures and Ratios
In supported planar tasks, the ratio of actuator to chamber pressure should be maximized. In self-supporting 3D tasks, a moderate chamber pressure should be selected to balance the tradeoff between the robot body being too soft to support itself against gravity and too stiff to steer.
In both classes of tasks, the pressures that can be chosen must have a chamber pressure above that required to grow at the desired speed [6], while both chamber and actuator pressures are below their respective burst pressures. Improved robot fabrication techniques that allow chamber and actuator pressures to increase seem promising for achieving both high growth speed and high curvature, especially with integrated pouch designs. Given current fabrication limitations, growth may need to pause temporarily while the robot steers into position, before reinflating the body chamber to resume extension. Control strategies that coordinate growth and steering – for example, alternating between “growth phases” and “steering phases” – could expand the range of environments where vine robots are deployed.
@@ VI Demonstration
The initial robot design representing the case where the design and control parameters were uninformed was inspired by the design presented in [3], which struggled to steer in the field while carrying its tip-mounted camera. That robot had a tip load of 145 g, maximum chamber and actuator pressures of 10.34 kPa, a body diameter of 8 cm, and exterior actuators. That tip load was the lowest amount that supported the camera design, whereas the pressure maximums were based on a similar design previously deployed in the field [4]. To mimic this robot, we used a design with an 8 cm body diameter and exterior actuators, with both the chamber and actuators having pressure maximums of 10.34 kPa. The tip mount was a plastic sleeve with 150 g attached to it for simplicity. When operating at its maximum chamber and actuator pressure, as was done in [3], it had a pressure ratio of 1. Under these conditions, it could only lift its tip over a 2 cm obstacle (Fig. 8(a)). For the growing task, the robot grew 21.3 cm, denoted by the yellow tape marks shown in Fig. 8, before curving in the plane. It required 10.34 kPa to grow. It was able to steer approximately 22° as estimated from the video (Fig. 8(b)).

### v2

@@ IV-C Experimental Setup
We conducted a series of planar bending trials with both body types (Fig. 5). The robots were held at a constant length of 1.75 m using a motorized spool within a vine robot base similar to the one in [CoadRAM2020]. Pressure was controlled using the same pressure regulators as in Sec. III-B. For each trial, a single actuator was aligned with the measurement plane, isolating its contribution to steering. Unlike the free-space trials in Sec. III, these experiments were performed with the robot lying on the floor, which reduced the influence of body weight, length, and gravitational loading along the body, allowing more direct characterization of fabrication methods and pressure ratio alone.
@@
The robot body was tracked using motion capture LED markers with the same system as in Sec. III-B, placed along the centerline at 25 cm spacing, starting at the base. This enabled reconstruction of the planar backbone for consistent extraction of the tip position and curvature. Similar to the previous experiments, we also placed LED markers aligned with the desired coordinate frame, this time with $y$ horizontal along the robot’s axis from base to tip when straight and $z$ up, and an LED at the base of the robot.
@@ V-B1 Diameter
Robot diameter primarily acts as a feasibility constraint rather than a continuous tuning parameter for steerability. In our experiments, the smallest tested diameter (3.2 cm) collapsed under a 50 g tip load at 6.89 kPa and a free length of 0.61 m, resulting in near-zero characteristic area (Fig. 3(d)). In contrast, all larger diameters (6.5 – 12.9 cm) remained stable and produced similar characteristic areas, horizontal ranges, vertical ranges, and initial vertical positions. These results indicate that once a vine robot’s diameter exceeds the minimum required to avoid collapse for a given load, pressure, and length, increasing diameter primarily rescales the resulting geometry rather than introducing new steering capability [blumenschein2018helical]. As a result, diameter selection should prioritize environmental constraints such as aperture size or available deployment volume, rather than steering performance.
@@ V-B2 Fabrication Method
The choice between exterior and integrated pouch fabrications determines both the pressure ratio required to initiate steering and the maximum curvature that can be achieved. As shown in Fig. 6, exterior pouch robots begin to curve at low pressure ratios ($P_{pouch}/P_{body}<1$), but their curvature saturates at higher ratios, limiting the maximum achievable bending. In contrast, integrated pouch robots exhibit no measurable curvature until the pressure ratio exceeds unity, after which curvature increases approximately linearly with pressure ratio and reaches higher peak values. These behaviors define two distinct operating regimes. Tasks that require steering while maintaining a high chamber pressure, such as growth under load or avoiding collapse in free space, benefit from exterior pouch fabrications, which can generate curvature at low ratios. Conversely, tasks performed in supported or quasi-static conditions, where higher pressure ratios are feasible, benefit from integrated pouch fabrications, which more effectively convert actuator pressure into large curvature once the hoop-stress threshold is exceeded.
@@ V-B3 Length
Robot length introduces a trade-off between lateral reach and vertical maneuverability. In our experiments, increasing free length consistently expanded the characteristic area and horizontal range, as a longer body amplifies tip displacement for a given curvature (Fig. 3(c)). At the same time, vertical range and initial vertical position decreased with length, indicating a growing influence of gravity that limits upward steering even before full collapse occurs.
As a result, increasing free length is beneficial for tasks dominated by horizontal steering, but can be detrimental for tasks that require lifting the tip or maintaining elevation in free space. While the robot’s length is inherently dictated by growth, an informed operator can exploit this trade-off through path planning. For example, when multiple apertures are available prior to a sharp turn, selecting an earlier aperture preserves greater usable length for lateral curvature while avoiding excessive gravitational penalty.
@@ V-B4 Tip Load
Tip load directly constrains free-space steerability by amplifying gravitational and transverse forces at the robot tip. In our experiments, increasing payload mass produced a nonlinear reduction in steerability, with characteristic area, horizontal range, and vertical range all degrading sharply beyond approximately 100 g (Fig. 3(a)). Above this load, the robot approached collapse and exhibited severely limited directional reach. These results indicate that tip load is not merely a performance modifier, but a dominant limiting factor for 3D steering. While minimizing payload mass is always beneficial, selectively removing or disabling non-essential sensors during challenging maneuvers offers a practical means of restoring steerability without redesigning the robot body.
@@ V-B5 Operating Pressures and Ratios
Operating pressures govern a fundamental trade-off between body stiffness, growth capability, and steering authority. In free-space 3D tasks, absolute chamber pressure exhibits a non-monotonic effect on steerability: low pressures provide insufficient support against gravity, while high pressures overly stiffen the body and limit actuator effectiveness, as reflected in both characteristic area and initial vertical position. In our experiments, characteristic area peaked at an intermediate chamber pressure, indicating an optimal stiffness range for unsupported steering. In supported planar tasks, steering performance is instead dominated by the relative pressure ratio between the pouch actuators and the body chamber. For both fabrication methods, curvature increased with $P_{pouch}/P_{body}$, but with distinct regimes: exterior pouch designs initiate steering at low ratios and saturate early, while integrated pouch designs require ratios exceeding unity before curving but achieve higher curvature at larger ratios.
Across both classes of tasks, feasible operating pressures are constrained by growth and structural limits. Chamber pressure must exceed the minimum required to sustain growth at the desired speed [HawkesScienceRobotics2017], while both chamber and actuator pressures must remain below their respective burst limits. Improved robot fabrication techniques that allow higher chamber and actuator pressures are particularly promising, as they would expand the achievable overlap between high growth speed and high curvature, especially for integrated pouch designs. Given current fabrication limitations, growth may need to pause temporarily while the robot steers into position, before reinflating the body chamber to resume extension. Control strategies that coordinate growth and steering – for example, alternating between “growth phases” and “steering phases” – could expand the range of environments where vine robots are deployed.
@@ VI Demonstration
The initial robot design representing the case where the design and control parameters were uninformed was inspired by the design presented in [mcfarlandSPROUT], which struggled to steer in the field while carrying its tip-mounted camera. That robot had a tip load of 145 g, maximum chamber and actuator pressures of 10.34 kPa, a body diameter of 8 cm, and exterior actuators. That tip load was the lowest amount that supported the camera design, whereas the pressure maximums were based on a similar design previously deployed in the field [CoadRAM2020]. To mimic this robot, we used a design with an 8 cm body diameter and exterior actuators, with both the chamber and actuators having pressure maximums of 10.34 kPa. The tip mount was a plastic sleeve with 150 g attached to it for simplicity. When operating at its maximum chamber and actuator pressure, as was done in [mcfarlandSPROUT], it had a pressure ratio of 1. Under these conditions, it could only lift its tip over a 2 cm obstacle (Fig. 8(a)). For the growing task, the robot grew 21.3 cm, denoted by the yellow tape marks shown in Fig. 8, before curving in the plane. It required 10.34 kPa to grow. It was able to steer approximately 22° as estimated from the video (Fig. 8(b)).